# Beyond Full Label: Single-Point Prompt for Infrared Small Target Label Generation

### v1

@@ Abstract
In this work, we make the first attempt to construct a learning-based single-point annotation paradigm for infrared small target label generation (IRSTLG). Our intuition is that label generation requires just one more point prompt than target detection: IRSTLG can be regarded as an infrared small target detection (IRSTD) task with the target location hint. Based on this insight, we introduce an energy double guided single-point prompt (EDGSP) [^1] Code will be made public at https://github.com/xdFai. framework, which adeptly transforms the target detection network into a refined label generation method. Specifically, the proposed EDGSP includes: 1) target energy initialization (TEI) to create a foundational outline for sufficient shape evolution of pseudo label, 2) double prompt embedding (DPE) for rapid localization of interested regions and reinforcement of individual differences to avoid label adhesion, and 3) bounding box-based matching (BBM) to eliminate false alarms. Experimental results show that pseudo labels generated by three baselines equipped with EDGSP achieve 100% object-level probability of detection (${{P}_{d}}$) and 0% false-alarm rate (${{F}_{a}}$) on SIRST, NUDT-SIRST, and IRSTD-1k datasets, with a pixel-level intersection over union (IoU) improvement of 13.28% over state-of-the-art label generation methods. Additionally, the downstream detection task reveals that our centroid-annotated pseudo labels surpass full labels, even with coarse single-point annotations, it still achieves 99.5% performance of full labeling.
@@ Introduction
Single-frame infrared small target detection (IRSTD) plays a crucial role in maritime rescue, early warning, and high-precision navigation (Wang, Zhou, and Wang 2019; Kou et al. 2023a, b; Zhang et al. 2024b). To isolate small targets from complex clutter, high-performance IRSTD methods typically extract discriminative features from extensive paired infrared images and pixel-level masks (Zhang et al. 2022; Wu et al. 2023; Chen et al. 2024a). This learning paradigm highlights the need for large-scale, accurate mask annotation. However, infrared (IR) small targets typically occupy only a few pixels, exhibit a low signal-to-clutter ratio, and lack well-defined edges. These characteristics render manual pixel-by-pixel annotation time-consuming and error-prone. Consequently, an efficient and reliable infrared small target label generation (IRSTLG) method is urgently required.
@@
Single-point annotation [^2] Unless explicitly stated, centroid single-point is used. (SA) has recently demonstrated efficiency in labeling IR small targets, reducing the time needed by 80% compared to pixel-level annotation. As shown in Figure 1(a), LESPS (Ying et al. 2023) made the first attempt to adopt SA for weakly supervised IRSTD. The mapping degradation is cleverly introduced to extract the target contour during network training. Then, these intermediate predictions are employed to guide the point label evolution. Precisely, LESPS integrates IRSTD network training with label generation into a unified framework. Subsequently, a standardized label generation paradigm emerged. As illustrated in Figure 1(c), these pioneering works use SA to guide traditional model optimization for generating pseudo masks, such as MCLC (Li et al. 2023a), MCGC (Kou et al. 2024), and COM (Li et al. 2024). Note that while these three works are claimed as “Single-Point Supervision”, they fundamentally belong to the label generation task with the single-point prompt as input. To avoid confusion, Figure 1(a)-(c) elucidates two critical links between these tasks. 1) If replace the full label in IRSTD with the pseudo label obtained in IRSTLG, single-point supervised IRSTD can be treated as an organic combination of single-point prompt IRSTLG and fully supervised IRSTD. This implies that developing IRSTLG can directly improve point-supervised IRSTD. 2) Label generation requires just one more point prompt than detection. Thus, IRSTLG can be regarded as an infrared small target detection task with target location hint. This indicates that the labeling task adheres to at least two essential criteria.
@@
To validate this reasoning, Figure 2 showcases a comparison between the detection results of five IRSTD methods (Wu et al. 2023; Hou et al. 2022; Li et al. 2022; Wu, Hong, and Chanussot 2022; Yuan et al. 2024) with the pseudo masks generated by three IRSTLG algorithms (Li et al. 2023a; Kou et al. 2024; Li et al. 2024). Experimental results indicate that existing IRSTLG methods fail to meet standards in AA and IoU. This issue arises because these traditional algorithms neglect to leverage target location cues to extract features from fully labeled data.
@@
Since well-designed IRSTD networks already achieve satisfactory target shape inference, and single-point prompts can be easily extracted from full labels, the primary goal of this paper is to design an effective “Upgrade Tool” that transforms the IRSTD network into a label generation method using the single-point prompt. As shown in Figure 1(d), the essence of this paradigm is learning point prompt-based mask labeling experience from already labeled data. In fact, point prompt combined with segmentation baseline has been applied to interactive annotation for general objects (Wang et al. 2019; Zhai et al. 2023). A common approach involves using multiple extreme points or corners to precisely crop the target region, then pairing it with the encoded prompt and full labels for fully supervised training. Although this pipeline has yielded satisfactory results in routine areas, directly applying it to our task reveals three pitfalls. 1) Missing edge coordinates. Only one pixel of the small target is annotated, the lack of edge coordinates hinders deep models focusing on the target contour. Even with multiple cropping scales and carefully crafted prompts, it still encounters obstacles in achieving accurate annotation (detailed see Experiments). 2) Unique target characteristics. Infrared small targets occupy only 0.15% of the total image pixels, and their edges are blurred by diffraction effects (Kou et al. 2023a). Two small targets in close proximity are easy to stick together and be mistakenly labeled as a single target, reducing ${{P}_{d}}$. 3) Strong background clutter. Cloud edges, sea phosphorescence, and bright buildings in IR images have similar energy distributions to small targets (Zhang et al. 2024a). Even under effective location cues, this similarity still causes the deep model to misclassify pixels in these interference regions, resulting in false alarms.
@@
Figure 2: Comparing IRSTD and IRSTLG methods on mixed SIRST, NUDT-SIRST, and IRSTD-1k in $IoU(\%)$, ${{P}_{d}}(\%)$, ${{F}_{a}}({10}^{-6})$. EDGSP-Coa. denotes EDGSP with the coarse single-point prompt.
@@
To tackle these problems, we propose a novel framework for IRSTLG, named energy double guided single-point prompt (EDGSP). Without requiring complex cropping, EDGSP enables end-to-end model target contours through the integration of local and global context provided by the point prompt and deep model. As shown in Figure 3, this framework consists of three key components. 1) Target energy initialization (TEI): Inspired by IR targets that always have similar energy and shape distribution, we expand each point prompt into a Gaussian with local salience priors as the initial pattern of every target to evolve more complete outlines. 2) Double prompt embedding (DPE): The encoded prompt is embedded twice. The first embedding ensures the model directly focuses on the interested region, even in shallow layers. The second embedding refines the target salience map, enhancing target boundaries and preventing shape merging. 3) Bounding box-based matching (BBM): As a post-processing approach, we generate bounding boxes for each candidate target, then match the point prompt with each box by coordinates to eliminate clutter.
@@
Benefiting from the above structures, as shown in Figure 2, our annotation method showcases powerful shape prediction capabilities and achieves ${{P}_{d}}$=100%, ${{F}_{a}}$=0on three datasets. The main contributions are summarised as follows.
@@
  • We propose the target energy initialization (TEI), double prompt embedding (DPE), and bounding box-based matching (BBM) strategies to address the insufficient label shape evolution, label adhesion, and false alarm.
@@
IRSTD techniques have been developed over decades, and for a long period, they relied on handcrafted traditional methods such as image filtering (Bai and Zhou 2010; Deshpande et al. 1999), human visual system (Kim and Lee 2012; Chen et al. 2013), and low-rank optimization (Gao et al. 2013; Liu et al. 2023). Recently, deep learning-based methods have been able to obtain high-level semantics from large amounts of paired data by stacking nonlinear feature extractors, achieving robust detection in complex scenes. Since pixel-by-pixel classification is crucial for subsequent target recognition, segmentation-based IRSTD is a key emphasis in these data-driven approaches. Based on a U-shaped structure (Ronneberger, Fischer, and Brox 2015), ACM (Dai et al. 2021) improves the feature fusion strategy of high-level semantics with low-level details using an asymmetric context modulator. To mitigate the semantic gap between encoders and decoders, DNA-Net (Li et al. 2022) adopts dense nested interactive modules to facilitate progressive interaction of cross-scale features and adaptive enhancement of fine-grained features, while UIU-Net (Wu, Hong, and Chanussot 2022) directly nest multiple small U-Net into each codec of U-shaped structure. To address the limitations in long-range information modeling, MTU-Net (Wu et al. 2023) employs a self-attention to compute the spatial correlations of fused features, thereby constructing a global representation of the image. Most recently, SCTransNet (Yuan et al. 2024) explicitly interacts and enhances full-level encoder features by spatial-channel cross transformer blocks to effectively suppress interference.
@@
One working mode related to us is point-supervised segmentation. As a weak supervision form, point supervision only requires one or a few points as label (Qian et al. 2019; Zhao and Yin 2020; Li et al. 2023b). Bearman et al. (Bearman et al. 2016) propose the first single-point supervised semantic segmentation method by integrating an objectiveness prior into a loss function, to elevate segmentation effectiveness. Subsequently, Zhang et al. (Zhang et al. 2023) leverage the common characteristics of target regions and present a contrast-based variational model for point-supervised histopathology image segmentation, achieving more regionally consistent and smoother boundaries. Close to our work, LESPS (Ying et al. 2023) is the first to introduce single-point supervision to infrared small target detection task, based on mapping degradation phenomenon to guide point label evolution.
Unlike the aforementioned works, we still rely on fully annotated masks for supervision, yet employ single-point as a form of input to guide the IRSTD network. This approach reveals how to utilize additional prompts to achieve high-precision pseudo-label generation for small targets.
@@
To the best of our knowledge, there are only three IRSTLG efforts. They are all based on the model-based single-point prompt methods. As a pioneer, MCLC (Li et al. 2023a) makes the first attempt to employ a Monte Carlo linear clustering process with random noise and averages the clustering results to obtain reliable pseudo masks. Subsequently, MCGC adaptive recovers small targets with multi-scale chain growth clustering, further enhancing annotation robustness (Kou et al. 2024). Recently, Li et al. (Li et al. 2024) designed a variational level set with an expected difference energy function to address the excessive regularization causing target contours to disappear. However, these works tend to treat the points as a ”Supervision” signal. This biased task orientation causes them to overlook the potential of combining point prompt-derived location cues to extract features from complete labeled datasets, culminating in pseudo labels with unclear edges, target adhesion, and false alarms.
In this paper, we accurately define the operating paradigm where images and single-point prompts serve as inputs and produce pseudo labels as “single-point prompt for infrared small target label generation”. Unlike previous methods, we present the first study of a learning-based paradigm for IRSTLG, which artfully evolves the target detection network into an advanced label generation method with the aid of the single-point prompt.
@@
The technique most similar to ours is point prompt interactive segmentation (Rother, Kolmogorov, and Blake 2004; Xu et al. 2016; Liew et al. 2017), which aims to reduce annotation time. DEXTR (Maninis et al. 2018) uses four extreme points for segmentation, the interest region is cropped and the Gaussian heat map is utilized as the encoding form of prompts. Similarly, IOG (Zhang et al. 2020) uses the approximate center of the target as the inside point and a set of symmetrical corner points as the outside points. After cropping the target of interest, the image, along with heat maps is fed into the segmentation model. Luo et al. (Luo et al. 2021) uses the exponentialized geodesic distance transform to convert medical images into Euclidean distance maps for improved foreground-background separation. Recently, SAM (Kirillov et al. 2023) has gained recognition for its flexibility and generalization capabilities. It is a transformer-based model for semantic segmentation that utilizes various types of prompts to encode key areas.
Despite the success of the above work on generic target segmentation, the unique characteristics and annotation patterns of infrared small targets render these strategies ineffective. The proposed EDGSP differs from them in three ways: 1) It processes the entire image directly without cropping. 2) It assigns explicit physical meaning to Gaussian heat maps, thereby using the double embedding strategy. 3) EDGSP introduces a box-based matching approach specifically designed to eliminate false alarms.
@@
The overview of our proposed EDGSP is shown in Figure 3. It is very concise and consists of three parts.
@@
The long-range imaging mechanism of infrared systems brings severe diffraction effects, resulting in small targets appearing with a nearly Gaussian distribution in infrared images (Dai et al. 2023). By combining this physical phenomenon with the local contrast prior of small targets (Chen et al. 2013), we propose target energy initialization to predefine a basic profile for each target. As shown in Figure 3, given an input image ${\mathbf{{I}_{i}}}\in\mathbb{R}^{{1}\times{H}\times{W}}$ and corresponding single-point prompt ${\mathbf{{I}_{p}}}\in\mathbb{R}^{{1}\times{H}\times{W}}$ (Since the single-point prompt is a pixel within the target, it can be derived from the already labeled ground truth ${\mathbf{{G}_{t}}}\in\mathbb{R}^{{1}\times{H}\times{W}}$). 2D Gaussian function is used to expand each point in ${\mathbf{{I}_{p}}}$ as an energy blob, thus creating a single init shape map ${\mathbf{{I}_{g}}}\in\mathbb{R}^{{1}\times{H}\times{W}}$. This expansion simplifies the task by shifting from shaping area targets with points to refining contours using Gaussian blobs, aligning the prior distribution of the targets, and aiding the model in learning complete shape representations.
@@
where ${\mathcal{G}^{Conv}_{w/o}}$ stands for $\mathcal{G}$ that without the last convolution layer. In the second embedding, ${\mathbf{{I}_{g}}}$ is treated as a semantic map that reinforces the target quantity, it is concatenated with the ${\mathbf{L}}$ along the channel dimension to refine the model’s final mapping and prevent close-range target sticking.
@@
where ${\mathbf{{O}_{s}}}\in\mathbb{R}^{{1}\times{H}\times{W}}$ represents the saliency map, ${Conv}$ denotes the last layer of $\mathcal{G}$, typically constructed with a $1\times 1$ or $3\times 3$ convolution that squeezes channel and preserves the spatial dimensions. The binary cross entropy loss is then used to compute pixel-wise differences between the ${\mathbf{{O}_{s}}}$ and the ground truth ${\mathbf{{G}_{t}}}$.
Table 1: Comparison with SOTA IRSTLG methods on SIRST, NUDT-SIRST, IRSTD-1k in terms of $IoU(\%)$, ${{P}_{d}}(\%)$, ${{F}_{at}}(\%)$, ${{F}_{a}}({10}^{-6})$. The best and the second-best results are highlighted in bold and underlined, respectively. EDGSP-Coa. indicates the EDGSP under the coarse single-point prompt.
@@
In previous works, eight connective regions matching is popular to eliminate false alarms during the label evolution process (Ying et al. 2023; Kou et al. 2024), ensuring that evolution information consistently applies to the initial point. However, this strategy ignores cases where the point prompt is located on or outside the target boundary, leading to incorrect processing results [^3] The shortcomings of this approach are detailed in the supplementary material.. We tackle this issue by adding a compact design. In Figure 3(c), after performing eight-neighborhood clustering on the threshold-segmented candidate map ${\mathbf{{O}_{c}}}$, we generate bounding boxes for each component to create a b-box map ${\mathbf{{B}_{m}}}$. Then, match the point prompt ${\mathbf{{I}_{p}}}$ with ${\mathbf{{B}_{m}}}$. Specifically, if the point in ${\mathbf{{I}_{p}}}$ falls inside a bounding box, the target in this box is retained; otherwise, it is discarded as a false positive. After performing this region match, we got the final pseudo label ${\mathbf{{O}_{p}}}$.
@@
Datasets. Experiments are conducted on three mixed datasets: SIRST (Dai et al. 2021), NUDT-SIRST (Li et al. 2022), and IRSTD-1k (Zhang et al. 2022), which include 427, 1327, and 1001 infrared images with one or more targets, respectively. We used the method from (Li et al. 2022) to partition the datasets for SIRST and NUDT-SIRST, and followed (Zhang et al. 2022) for splitting IRSTD-1k.
@@
Implement Details. SCTransNet (Yuan et al. 2024) is used as the backbone of EDGSP. The peak half-width of the Gaussian is set to 4 and the segmentation threshold is set to 0.5. We compare all IRSTLG methods: MCLC (Li et al. 2023a), MCGC (Kou et al. 2024) and COM (Li et al. 2024), and learning-based single-point annotation methods: SAM (Kirillov et al. 2023) and Robust SAM (Chen et al. 2024b). Furthermore, we also include the point-supervised IRSTD framework LESPS (Ying et al. 2023), using DNA-Net as the backbone, consistent with the original paper. We provide more details of experiments in the supplementary material.
Table 2: The improvement brought by the EDGSP framework to seven SOTA IRSTD methods on SIRST, NUDT-SIRST, IRSTD-1k in terms of $IoU(\%)$, ${{P}_{d}}(\%)$, ${{F}_{at}}(\%)$, ${{F}_{a}}({10}^{-6})$, $Flops(G)$, and $Params(M)$.
@@
As shown in Table 1, only EDGSP achieves ${{P}_{d}}$=100%, ${{F}_{a}}$=0%, ${{F}_{at}}$=0across three datasets at the same time. This indicates that our framework is currently the only one ensuring the most basic level of annotation quality. Additionally, EDGSP’s IoU surpasses the state-of-the-art (SOTA) labeling method (MCGC) by 14.27, demonstrating a superior shape prediction capability. Even with the coarse single-point prompt, EDGSP outperforms other methods with the centroid single-point prompt.

### v4

@@ Abstract
In this work, we make the first attempt to construct a learning-based single-point annotation paradigm for infrared small target label generation (IRSTLG). Our intuition is that label generation requires just one more point prompt than target detection: IRSTLG can be regarded as an infrared small target detection (IRSTD) task with the target location hint. Based on this insight, we introduce an energy double guided single-point prompt (EDGSP) [^1] The code is available at https://github.com/xdFai/EDGSP. framework, which adeptly transforms the target detection network into a refined label generation method. Specifically, the proposed EDGSP includes: 1) target energy initialization (TEI) to create a foundational outline for sufficient shape evolution of pseudo label, 2) double prompt embedding (DPE) for rapid localization of interested regions and reinforcement of individual differences to avoid label adhesion, and 3) bounding box-based matching (BBM) to eliminate false alarms. Experimental results show that pseudo labels generated by three baselines equipped with EDGSP achieve 100% object-level probability of detection (${{P}_{d}}$) and 0% false-alarm rate (${{F}_{a}}$) on SIRST, NUDT-SIRST, and IRSTD-1k datasets, with a pixel-level intersection over union (IoU) improvement of 13.28% over state-of-the-art (SOTA) label generation methods. In the practical application of downstream IRSTD, EDGSP realizes, for the first time, a single-point generated pseudo mask beyond the full label. Even with coarse single-point annotations, it still achieves 99.5% performance of full labeling.
@@ Introduction
Single-frame infrared small target detection (IRSTD) plays a crucial role in maritime rescue, early warning, and high-precision navigation (Wang, Zhou, and Wang 2019; Kou et al. 2023a, b; Zhang et al. 2024b). To isolate small targets from clutter, high-performance IRSTD methods typically extract discriminative features from extensive paired infrared images and pixel-level masks (Zhang et al. 2022; Wu et al. 2023; Chen et al. 2024a). This learning paradigm highlights the need for large-scale, accurate mask annotations. However, infrared (IR) small targets typically occupy only a few pixels and exhibit a low signal-to-clutter ratio and poorly defined edges. These characteristics render manual pixel-by-pixel annotation time-consuming and error-prone. Consequently, an efficient and reliable infrared small target label generation (IRSTLG) method is urgently required.
@@
Single-point annotation [^2] Unless explicitly stated, centroid single-point is used. (SA) has recently demonstrated efficiency in labeling IR small targets, reducing the time needed by 80% compared to pixel-level annotation (Li et al. 2023a). As shown in Figure 1(a), LESPS (Ying et al. 2023) made the first attempt to adopt SA for weakly supervised IRSTD. The mapping degradation is cleverly introduced to extract the target contour during network training. Then, these intermediate predictions are employed to guide the point label evolution. Precisely, LESPS integrates IRSTD network training with label generation into a unified framework. Subsequently, a standardized label generation paradigm emerged. As illustrated in Figure 1(c), these pioneering works use SA to guide traditional model optimization for generating pseudo labels, such as MCLC (Li et al. 2023a), MCGC (Kou et al. 2024), and COM (Li et al. 2024). Note that while these three works are claimed as “single-point supervision”, they fundamentally belong to the label generation task with the “single-point prompt” as input. To avoid confusion, Figure 1(a)-(c) elucidates two critical links between these tasks. 1) If replace the full label in IRSTD with the pseudo label obtained in IRSTLG, single-point supervised IRSTD can be treated as an organic combination of single-point prompt IRSTLG and fully supervised IRSTD. This implies that developing IRSTLG can directly improve point-supervised IRSTD. 2) Label generation requires just one more point prompt than detection. Thus, IRSTLG can be regarded as an infrared small target detection task with target location hint. This indicates that the labeling task adheres to at least two essential criteria.
@@
To validate this reasoning, Figure 2 showcases a comparison between the detection results of five IRSTD methods (Wu et al. 2023; Hou et al. 2022; Li et al. 2022; Wu, Hong, and Chanussot 2022; Yuan et al. 2024) with the pseudo masks generated by three IRSTLG algorithms (Li et al. 2023a; Kou et al. 2024; Li et al. 2024). Experimental results indicate that existing IRSTLG methods fail to meet standards in AA and IoU, primarily because these traditional algorithms neglect to leverage target location cues for feature extraction from fully labeled data.
@@
Since well-designed IRSTD networks already achieve satisfactory target shape inference, and single points can be easily extracted from full labels, the primary goal of this paper is to design an effective “upgrade tool” that transforms the IRSTD network into a more advanced label generation method under the single-point prompt. As shown in Figure 1(d), the essence of this paradigm is learning point prompt-based mask labeling experience from already labeled data. In fact, point prompt combined with segmentation baseline has been applied to interactive annotation for general objects (Wang et al. 2019; Zhai et al. 2023). A common approach involves using multiple extreme points or corners to precisely crop the target region, then pairing it with the encoded prompt and full labels for fully supervised training. Although this pipeline has yielded satisfactory results in routine areas, directly applying it to our task exposes three pitfalls. 1) Missing edge coordinates. Only one pixel of the small target is annotated, the lack of edge coordinates hinders deep models focusing on the target contour. Even with multiple cropping scales and carefully crafted prompts, it still encounters obstacles in achieving accurate annotation (detailed see Experiments). 2) Unique target characteristics. Infrared small targets occupy only 0.15% of the total image pixels, and their edges are blurred by diffraction effects (Kou et al. 2023a). Two small targets in close proximity can easily stick together and be mistakenly labeled as a single target, reducing ${{P}_{d}}$. 3) Strong background clutter. Cloud edges, sea phosphorescence, and bright buildings in IR images have similar energy distributions to small targets (Zhang et al. 2024a). Even with valid center-coordinate cues, this similarity still causes the deep model to misclassify pixels in interference regions, resulting in false alarms.
@@
Figure 2: Comparing IRSTD and IRSTLG methods on mixed datasets from SIRST, NUDT-SIRST, and IRSTD-1k in $IoU(\%)$, ${{P}_{d}}(\%)$, and ${{F}_{a}}({10}^{-6})$. EDGSP-Coa. denotes EDGSP with the coarse single-point prompt.
@@
To tackle these problems, we propose a novel framework for IRSTLG, named energy double guided single-point prompt (EDGSP). Without requiring complex cropping, EDGSP enables end-to-end modeling target contours through the integration of local and global context provided by the point prompt and deep network. As shown in Figure 3, this framework consists of three components. 1) Target energy initialization (TEI): Inspired by IR targets that always have similar energy and shape distribution, we expand each point prompt into a Gaussian with local salience priors as the initial pattern of every target to evolve more complete outlines. 2) Double prompt embedding (DPE): The encoded prompt is embedded twice. The first embedding ensures that the model directly focuses on the region of interest, even in shallow layers. The second embedding refines the target salience map, fine-tuning mask boundaries and preventing shape merging. 3) Bounding box-based matching (BBM): As a post-processing step, we generate bounding boxes for each candidate target and then match the point prompt with each box by coordinates to eliminate clutter.
@@
Benefiting from the above structures, as shown in Figure 2, our method showcases powerful shape prediction capabilities and achieves ${{P}_{d}}$=100%, ${{F}_{a}}$=0on three datasets. The main contributions are summarised as follows.
@@
  • We propose target energy initialization (TEI), double prompt embedding (DPE), and bounding box-based matching (BBM) strategies to address insufficient shape evolution, label adhesion, and false alarms.
@@
IRSTD techniques have been developed over decades, and for a long period, they relied on handcrafted traditional methods such as image filtering (Bai and Zhou 2010; Deshpande et al. 1999), human visual system (Kim and Lee 2012; Chen et al. 2013), and low-rank optimization (Gao et al. 2013; Liu et al. 2023). Recently, deep learning-based methods have been able to obtain high-level semantics from large amounts of paired data by stacking nonlinear feature extractors, achieving robust detection in complex scenes. Since pixel-by-pixel classification is crucial for subsequent target recognition, segmentation-based IRSTD is a key emphasis in these data-driven approaches. Based on a U-shaped structure (Ronneberger, Fischer, and Brox 2015), ACM (Dai et al. 2021) improves the feature fusion strategy of high-level semantics with low-level details using an asymmetric context modulator. To mitigate the semantic gap between encoders and decoders, DNA-Net (Li et al. 2022) adopts dense nested interactive modules to facilitate progressive interaction of cross-scale features and adaptive enhancement of fine-grained features, while UIU-Net (Wu, Hong, and Chanussot 2022) directly nests the small U-Net into each codec of the large U-shaped structure. To address the limitations in long-range information modeling, MTU-Net (Wu et al. 2023) employs a self-attention to compute the spatial correlations of fused features, thereby constructing a global representation of the image. Most recently, SCTransNet (Yuan et al. 2024) explicitly interacts and enhances full-level encoder features by spatial-channel cross transformer blocks to effectively suppress interference.
@@
One working mode related to us is point-supervised segmentation. As a manner of weak supervision, point supervision only requires one or a few points as labels (Qian et al. 2019; Zhao and Yin 2020; Li et al. 2023b). Bearman et al. (Bearman et al. 2016) propose the first single-point supervised semantic segmentation method by integrating an objectiveness prior into a loss function, to elevate segmentation effectiveness. Subsequently, Zhang et al. (Zhang et al. 2023) leverage the common characteristics of target regions and present a contrast-based variational model for point-supervised histopathology image segmentation, achieving more regionally consistent and smoother boundaries. Close to our work, LESPS (Ying et al. 2023) is the first to introduce single-point supervision to the infrared small target detection task, using the mapping degradation phenomenon to guide point label evolution.
Unlike the aforementioned works, we still rely on fully annotated masks for supervision, yet employ single-point as a form of input to guide the IRSTD network. This approach investigates how to utilize the additional prompt to achieve high-precision pseudo-label generation for IR small targets.
@@
To the best of our knowledge, there are only three IRSTLG efforts. They are all model-based single-point prompt methods. As a pioneer, MCLC (Li et al. 2023a) makes the first attempt to employ a Monte Carlo linear clustering process with random noise and averages the clustering results to obtain reliable pseudo masks. Subsequently, MCGC adaptive recovers small targets with multi-scale chain growth clustering, further enhancing annotation robustness (Kou et al. 2024). Li et al. (Li et al. 2024) design a variational level set with an expected difference energy function to address the excessive regularization causing target contours to disappear. However, these works tend to treat the point as a ”Supervision” signal. This biased task orientation causes them to overlook the potential of combining point prompt-derived location cues to extract features from complete labeled datasets, culminating in pseudo labels with unclear edges, target adhesion, and false alarms.
In this paper, we accurately define the operating paradigm where images and single-point prompts serve as inputs and produce pseudo labels as “single-point prompt for infrared small target label generation”. Unlike previous methods, we present the first study of a learning-based paradigm for IRSTLG, which artfully evolves the target detection network into a high-level label generation method with the aid of the single-point prompt.
@@
The technique most similar to ours is point prompt interactive segmentation (Rother, Kolmogorov, and Blake 2004; Xu et al. 2016; Liew et al. 2017), which aims to reduce annotation time. DEXTR (Maninis et al. 2018) uses four extreme points for segmentation, the interest region is cropped and the Gaussian heatmap is utilized as the encoding form of prompts. Similarly, IOG (Zhang et al. 2020) uses the approximate center of the target as the inside point and a set of symmetrical corner points as outside points. After cropping the target of interest, the image, along with the heatmap is fed into the segmentation model. Luo et al. (Luo et al. 2021) uses the exponentialized geodesic distance transform to convert medical images into Euclidean distance maps for improved foreground-background separation. Recently, SAM (Kirillov et al. 2023) has gained recognition for its flexibility and generalization capabilities. It is a transformer-based model for semantic segmentation that utilizes various types of prompts to encode key areas.
Despite the success of the above work on generic target segmentation, the unique characteristics and annotation patterns of infrared small targets render these strategies ineffective. The proposed EDGSP differs from them in three ways: 1) It processes the entire image directly without cropping. 2) It assigns explicit physical meaning to Gaussian heatmap and introduces the double embedding strategy. 3) EDGSP adopts a box-based matching approach customized to eliminate false alarms.
@@
The overview of our proposed EDGSP is shown in Figure 3. It is concise and consists of three parts.
@@
The long-range imaging mechanism of infrared systems brings severe diffraction effects, resulting in small targets appearing with a nearly Gaussian distribution in infrared images (Dai et al. 2023). By combining this physical phenomenon with the local contrast prior of small targets (Chen et al. 2013), we propose target energy initialization to predefine a basic profile for each target. As shown in Figure 3, given an input IR image ${\mathbf{{I}_{i}}}\in\mathbb{R}^{{1}\times{H}\times{W}}$ and corresponding single-point prompt ${\mathbf{{I}_{p}}}\in\mathbb{R}^{{1}\times{H}\times{W}}$ (Since the single-point prompt is a pixel within the target, it can be derived from the already labeled ground truth ${\mathbf{{G}_{t}}}\in\mathbb{R}^{{1}\times{H}\times{W}}$). 2D Gaussian function is used to expand each point in ${\mathbf{{I}_{p}}}$ as an energy blob, thus creating a single init shape map ${\mathbf{{I}_{g}}}\in\mathbb{R}^{{1}\times{H}\times{W}}$. This expansion simplifies the task by shifting from shaping area targets with points to refining contours using Gaussian blobs, aligning the prior distribution of the object, and aiding the model in learning complete shape representations.
@@
where ${\mathcal{G}^{Conv}_{w/o}}$ stands for $\mathcal{G}$ that without the last convolution layer. In the second embedding, ${\mathbf{{I}_{g}}}$ is treated as a semantic map that emphasizes the target quantity and edge difference. It is concatenated with the ${\mathbf{L}}$ along the channel dimension to refine the model’s final mapping and prevent close-range target sticking.
@@
where ${\mathbf{{O}_{s}}}\in\mathbb{R}^{{1}\times{H}\times{W}}$ represents the saliency map, ${Conv}$ denotes the last layer of $\mathcal{G}$, typically constructed with a $1\times 1$ or $3\times 3$ convolution that squeezes channel and preserves the spatial dimensions. The binary cross entropy loss is then employed to compute pixel-wise differences between the ${\mathbf{{O}_{s}}}$ and the ground truth ${\mathbf{{G}_{t}}}$.
Table 1: Comparison with SOTA IRSTLG methods on SIRST, NUDT-SIRST, and IRSTD-1k in terms of $IoU(\%)$, ${{P}_{d}}(\%)$, ${{F}_{at}}(\%)$, and ${{F}_{a}}({10}^{-6})$. The best and the second-best results are highlighted in bold and underlined, respectively. EDGSP-Coa. indicates the EDGSP under the coarse single-point prompt.
@@
In previous works, eight connective regions matching is commonly used to exclude false alarms during the label evolution process (Ying et al. 2023; Kou et al. 2024), ensuring that evolution information consistently applies to the initial point. However, this strategy ignores cases where the point prompt is located on or outside the target boundary, leading to incorrect processing results [^3] The shortcomings of eight connective regions matching are detailed in the supplementary material.. We tackle this issue by adding a compact design. In Figure 3(c), after performing eight neighborhood clustering on the threshold-segmented candidate map ${\mathbf{{O}_{c}}}$, we generate bounding boxes for each component to create a b-box map ${\mathbf{{B}_{m}}}$. Then, match the point prompt ${\mathbf{{I}_{p}}}$ with ${\mathbf{{B}_{m}}}$. Specifically, if the point in ${\mathbf{{I}_{p}}}$ falls inside a bounding box, the target in this box is retained; otherwise, it is discarded as a false positive. After performing this region match, we got the final pseudo label ${\mathbf{{O}_{p}}}$.
@@
Datasets. Following (Ying et al. 2023), experiments are conducted on three mixed datasets: SIRST (Dai et al. 2021), NUDT-SIRST (Li et al. 2022), and IRSTD-1k (Zhang et al. 2022), which include 427, 1327, and 1001 infrared images with one or more targets, respectively. We use (Li et al. 2022) to partition the datasets for SIRST and NUDT-SIRST, and follow (Zhang et al. 2022) to split IRSTD-1k.
@@
Implement Details. SCTransNet (Yuan et al. 2024) is used as the backbone of EDGSP. The peak half-width of the Gaussian is set to 4, and the segmentation threshold is set to 0.5. We compare all IRSTLG methods: MCLC (Li et al. 2023a), MCGC (Kou et al. 2024) and COM (Li et al. 2024), and learning-based single-point annotation methods: SAM (Kirillov et al. 2023) and Robust SAM (Chen et al. 2024b). Furthermore, we also include the point-supervised IRSTD framework LESPS (Ying et al. 2023), using DNA-Net as the backbone, consistent with the original paper. More details about the experiments are provided in the supplementary material.
Table 2: The improvement brought by the EDGSP framework to seven SOTA IRSTD methods on mixed datasets from SIRST, NUDT-SIRST, and IRSTD-1k in terms of $IoU(\%)$, ${{P}_{d}}(\%)$, ${{F}_{at}}(\%)$, ${{F}_{a}}({10}^{-6})$, $Flops(G)$, and $Params(M)$.
@@
As shown in Table 1, only EDGSP achieves ${{P}_{d}}$=100%, ${{F}_{a}}$=0%, ${{F}_{at}}$=0across three datasets at the same time. This indicates that our framework is currently the only one guaranteeing the most basic level of annotation quality. Additionally, EDGSP’s IoU surpasses the state-of-the-art (SOTA) labeling method (MCGC) by 14.27, demonstrating a superior shape prediction capability. Even with the coarse single-point prompt, EDGSP still outperforms other methods that use the centroid single-point prompt.